suppose we have two investment ptojects , both with initial cost of one billion.Both projects have zero return in period year 1, when they are built. project 1 return 0 in period 1 and 4 in period 2. project two on the other hand returns 2 million in period 1 and 1 million in period 2. the costs and returns of two projects are summarized in the table below?

1
Expert's answer
2022-03-21T12:48:51-0400

Solution

"NPV = \frac{F}{(1 + r)^n }"

Where, PV = Present Value,

F = Future payment (cash flow),

r = Discount rate,

n = the number of periods in the future profiles of the two different projects and observe their point of intersection.


Project 1

NPV = Today's value of the expected cash flows − Today's value of invested cash.


"NPV=4,000,000,000-1,000,000,000"

"NPV=3,000,000,000"



Project 2

NPV = Today's value of the expected cash flows − Today's value of invested cash.


"NPV=2,000,000,000-1,000,000,000"

"NPV=1,000,000,000"


Net present value, or NPV, is used to calculate the current total value of a future stream of payments. If the NPV of a project or investment is positive, it means that the discounted present value of all future cash flows related to that project or investment will be positive, and therefore attractive. Therefore, project 1 is viable because the NPV is higher as compared to project 2.